OpenStudy (anonymous):

solve for x. show all steps leading to the solution 3(2-x)=5(2x-7)+2

OpenStudy (anonymous):

3(2-x)=5(2x-7)+2 6-3x = 10x -35 +2 6 -3x = 10x -33 39 = 13x 39/13 =x 3 =x

OpenStudy (anonymous):

6-3x=10x-35+2 6+33=10+3x 39=13x x=39/13 x=3